# Randomization-based Inference for MCP-Mod

### v1

@@ 5.2 Future Research Directions
The efficacy of randomization tests could be further examined using different test statistics. We evaluated test statistics motivated by the MCP-Mod approach, but also other test statistics could be employed offering improved performance under specific conditions.
Given the promising results, future research could explore the broader application of randomization tests in pharmacometric models and trial designs with additional adaptive elements. These trials often involve complex designs and high-dimensional data, where randomization-based approaches could offer robust solutions, especially in the presence of small sample sizes, unbalanced treatment groups and time trends.

### v2

@@ 5.2 Future Research Directions
To address the issues related to complete separation, we chose Firth’s method. However, alternative approaches could be explored, such as adding one success and one failure to each arm, as suggested by Clogg [39].
@@
The efficacy of randomization tests could be further examined using different test statistics. We evaluated test statistics motivated by the MCP-Mod approach, but also other test statistics could be employed offering improved performance under specific conditions. Notably, we also did not alter the modeling part of the MCP-Mod framework. Exploring randomization based inference for other dose-finding methods would also be of interest.
Given the promising results, future research could explore the broader application of randomization tests in pharmacometric models and trial designs with additional adaptive elements. These trials often involve complex designs and high-dimensional data, where randomization-based approaches could offer robust solutions, especially in the presence of small sample sizes, unbalanced treatment groups and time trends. A natural extension would be extending our work to the longitudinal MCP-Mod framework by Buatois et al. (2021) [12].